# Ctrl-X: Controlling Structure and Appearance for Text-To-Image Generation Without Guidance

### v1

@@ Abstract
Recent controllable generation approaches such as FreeControl [24] and Diffusion Self-guidance [7] bring fine-grained spatial and appearance control to text-to-image (T2I) diffusion models without training auxiliary modules. However, these methods optimize the latent embedding for each type of score function with longer diffusion steps, making the generation process time-consuming and limiting their flexibility and use. This work presents Ctrl-X, a simple framework for T2I diffusion controlling structure and appearance without additional training or guidance. Ctrl-X designs feed-forward structure control to enable the structure alignment with a structure image and semantic-aware appearance transfer to facilitate the appearance transfer from a user-input image. Extensive qualitative and quantitative experiments illustrate the superior performance of Ctrl-X on various condition inputs and model checkpoints. In particular, Ctrl-X supports novel structure and appearance control with arbitrary condition images of any modality, exhibits superior image quality and appearance transfer compared to existing works, and provides instant plug-and-play functionality to any T2I and text-to-video (T2V) diffusion model. See our project page for an overview of the results: https://genforce.github.io/ctrl-x.
00footnotetext: *Equal contribution.
@@ 1 Introduction
To tackle these limitations, we present Ctrl-X, a simple training-free and guidance-free framework for T2I diffusion with structure and appearance control. We name our method “Ctrl-X” because we reformulate the controllable generation problem by ‘cutting’ (and ‘pasting’) two tasks together: Spatial structure preservation and semantic-aware stylization. Our insight is that diffusion feature maps capture rich spatial structure and high-level appearance from early diffusion steps sufficient for structure and appearance control without guidance. To this end, Ctrl-X employs feature injection and spatially-aware normalization in the attention layers to facilitate structure and appearance alignment with user-provided images. By being guidance-free, Ctrl-X eliminates additional optimization overhead and sampling steps, resulting in a $40$-fold increase in inference speed compared to guidance-based methods. Figure 1 shows some generation results. Moreover, Ctrl-X supports arbitrary structure conditions beyond natural images and can be applied to any T2I and even text-to-video (T2V) diffusion models. Extensive quantitative and qualitative experiments demonstrate the superior image quality and appearance alignment of our method over prior works.
@@ 4.1 Feed-forward structure control
In contrast to [36, 18, 24], we do not perform inversion and instead directly use forward diffusion (Equation 1) to obtain $\mathbf{x}^{\mathrm{s}}_{t}$. We observe that $\mathbf{x}^{\mathrm{s}}_{t}$ obtained via the forward diffusion process contains sufficient structure information even at very early/high time steps, as shown in Figure 3. This also reduces appearance leakage common to inversion-based methods observed by FreeControl [24]. We study our feed-forward structure control method in Sections 5.1 and 5.3.
@@ 5.1 T2I diffusion with structure and appearance control
Baselines.    For training-based methods, ControlNet [44] and T2I-Adapter [25] learn an auxiliary module that injects a condition image into a pretrained diffusion model for structure alignment. We then combine them with IP-Adapter [43], a trained module for image prompting and thus appearance transfer. Splicing ViT Features [35] trains a U-Net from scratch per source-appearance image pair to minimize their DINO-ViT self-similarity distance and global [CLS] token loss. (For structure conditions not supported by a training-based baseline, we convert them to canny maps.) For guidance-based methods, FreeControl [24] enforce structure and appearance alignment via backpropagated score functions computed from diffusion feature subspaces. For guidance-free methods, Cross-Image Attention [1] manipulates attention weights to transfer appearance while maintaining structure. We run all methods on SDXL v1.0 [27] when possible and on their default base models otherwise.
@@
Evaluation metrics.    For quantitative evaluation, we report two widely-adopted metrics: DINO Self-sim measures the self-similarity distance [35] between the structure and output image in the DINO-ViT [6] feature space, where a lower distance indicates better structure preservation; DINO CLS measures the loss between the DINO-ViT global [CLS] tokens of the appearance and output image [35], where a lower loss indicates better appearance alignment.
@@
Qualitative results.    As shown in Figures 4 and 5, Ctrl-X faithfully preserves structure from structure images ranging from natural images and ControlNet-supported conditions (e.g. HED, segmentation) to in-the-wild conditions (e.g. wireframe, 3D mesh) not possible in prior training-based methods while adeptly transferring appearance from the appearance image with semantic correspondence.
@@
Comparison to baselines.    Figure 5 and Table 1 compare Ctrl-X to the baselines. For training-based and guidance-based methods, despite T2I-Adapter [25] and FreeControl’s [24] stronger structure preservation (smaller DINO self-similarity distances), they generally struggle to enforce faithful appearance transfer and yield worse global CLS losses, which is particularly visible in Figure 5 row 1 and 3. Since the training-based methods combine a structure control module (ControlNet [44] and T2I-Adapter) with a separately-trained appearance transfer module IP-Adapter [43], the two modules sometimes exert conflicting control signals at the cost of appearance transfer (e.g. row 1)—and for ControlNet, structure preservation as well. For FreeControl, its appearance score function from extracted embeddings may not sufficiently capture more complex appearance correspondences, which, along with needing per-image hyperparameter tuning, results in lower contrast outputs and sometimes failed appearance transfer (e.g. row 4). Moreover, despite Splicing ViT Features [35] having the best DINO self-similarity and CLS scores in Table 1, Figure 5 reveals that its output images are often blurry while displaying structure image appearance leakage with non-natural images (e.g. row 3, 5, and 6). It benchmarks well because its per-image training minimizes these two metrics directly.
@@
Guidance-free baseline Cross-Image Attention [1], in contrast, is less robust and more sensitive to the structure image’s appearance, as the inverted structure latents contain strong appearance information. This causes both poorer structure alignment and frequent appearance leakage or artifacts (e.g. row 6) from the structure to the output images, resulting in worse DINO self-similarity distances and global CLS losses. In practice, we find Cross-Image Attention to be sensitive to its masking domain and sometimes fails to produce outputs with crossmodal pairs (e.g. wireframes to photos).
Table 1: Quantitative comparison of structure and appearance control.    Ctrl-X consistently outperforms both training-based and training-free methods in appearance alignment and shows comparable or better structure preservation compared to training-based and guidance-free methods, measured by DINO ViT self-similarity and global CLS token loss [35], respectively.
Method Training Natural image ControlNet-supported New condition Inference time (s)
Self-sim $\downarrow$ DINO CLS $\downarrow$ Self-sim $\downarrow$ DINO CLS $\downarrow$ Self-sim $\downarrow$ DINO CLS $\downarrow$
Splicing ViT Features [35] 0.030 0.006 0.043 0.012 0.037 0.013 4289.20
ControlNet + IP-Adapter [44, 43] 0.068 0.109 0.136 0.092 0.139 0.103 23.10
T2I-Adapter + IP-Adapter [25, 43] 0.055 0.119 0.118 0.118 0.109 0.131 17.70
Cross-Image Attention [1] 0.145 0.110 0.196 0.152 0.195 0.139 216.46
FreeControl [24] 0.058 0.132 0.101 0.119 0.089 0.139 1210.02
Ctrl-X (ours) 0.057 0.096 0.121 0.084 0.109 0.097 30.65
Figure 6: Qualitative comparison of conditional generation.    Ctrl-X displays comparable structure control and superior prompt alignment to training-based methods, and it also has better image quality and is more robust than guidance-based and -free methods across different conditions.
@@
Inference efficiency.    We study the inference time of our method compared to the baselines, all with base model SDXL v1.0 except Cross-Image Attention (SD v1.5) and Splicing ViT Features (U-Net). Table 1 reports the average inference time using a single NVIDIA A6000 GPU. Ctrl-X is slightly slower than training-based ControlNet (1.32×1.32$\times$1.32 ×) and T2I-Adapter (1.73×1.73$\times$1.73 ×) with IP-Adapter, yet it is significantly faster than per-image-trained Splicing ViT (0.0071×0.0071$\times$0.0071 ×), guidance-based FreeControl (0.025×0.025$\times$0.025 ×), and guidance-free Cross-Image Attention (0.14×0.14$\times$0.14 ×). Our training-free and guidance-free method achieves comparable runtimes to training-based methods, indicating its flexibility.
@@
Extension to prompt-driven conditional generation.    Ctrl-X also supports prompt-driven conditional generation, where it generates an output image complying with the given text prompt while aligning with the structure from the structure image, as shown in Figures 4 and 6. Inspired by FreeControl [24], instead of a given $\mathbf{I}^{\mathrm{a}}$, Ctrl-X can jointly generate $\mathbf{I}^{\mathrm{a}}$ based on the text prompt alongside $\mathbf{I}^{\mathrm{o}}$, where we obtain $\mathbf{x}^{\mathrm{a}}_{t-1}$ via denoising with Equation 2 from $\mathbf{x}^{\mathrm{a}}_{t}$ without control. Baselines, qualitative and quantitative analysis, and implementation details are available in Appendix B.
5.2 Extension to video diffusion models
Figure 7: Extension to text-to-video (T2V) models.    Ctrl-X can be directly applied to T2V models [9, 32] for controllable video structure and appearance control. More results are in Appendix C.
@@ 5.2 Extension to video diffusion models
Ctrl-X is training-free, guidance-free, and demonstrates competitive runtime. Thus we can directly apply our method to text-to-video (T2V) models, as seen in Figure 7. Our method closely aligns the structure between the structure and output videos while transferring temporally consistent appearance from the appearance image. Additional T2V results on more models can be found in Appendix C.
@@ 5.3 Ablations
Figure 8: Ablations.    We study ablations on (a) structure and appearance control, (b) appearance transfer methods, and (c) inversion vs. our method for obtaining noisy structure and output latents.
@@
Appearance transfer method.    As we consider appearance transfer as a stylization task, we compare our appearance statistics transfer with and without attention weighting. Without attention weighting (equivalent to AdaIN [15]), the normalization is global and thus cannot consider the semantic correspondence between the appearance and output images, so the outputs look low-contrast.
@@ 6 Conclusion
We present Ctrl-X, a training-free and guidance-free framework for structure and appearance control of any T2I and T2V diffusion model. Ctrl-X utilizes pretrained T2I diffusion model feature correspondences, supports arbitrary structure image conditions, works with multiple model architectures, and achieves competitive structure preservation and superior appearance transfer compared to training- and guidance-based methods while enjoying the low overhead benefits of guidance-free methods.
@@
Limitations.    As shown in Figure 9, the key limitation of Ctrl-X is the semantic-award appearance transfer method may fail to capture the target appearance when the instance is small because of the low resolution of the feature map. We hope our method and findings can unveil new possibilities and research on controllable generation as generative models become bigger and more capable.
@@
Appendix A Method and implementation details
Method Structure control Appearance control Training-free Guidance-free
Natural images In-the-wild conditions
ControlNet [44] (+ IP-Adapter [43])
T2I-Adapter [25] (+ IP-Adapter [43])
SDEdit [23]
Prompt2Prompt [11]
Plug-and-Play [36]
InfEdit [41]
Splicing ViT Attention [35]
Cross-Image Attention [1]
FreeControl [24]
Ctrl-X (ours)
We compare Ctrl-X to prior works in terms of capabilities in Table 2. Compared to baselines, our method is the only work which supports appearance and structure control with any structure conditions, while being training-free and guidance-free.
@@ Appendix A Method and implementation details
$n^{\mathrm{r}}$ times for (normalized) time steps $t\in[\tau^{\mathrm{r}}_{0},\tau^{\mathrm{r}}_{1}]$, where $\tau^{\mathrm{r}}_{0},\tau^{\mathrm{r}}_{1}\in[0,1]$.
@@
Experiment hyperparameters.    For both T2I diffusion with structure and appearance control and structure-only conditional generation, we use Stable Diffusion XL (SDXL) v1.0 [27] for all Ctrl-X experiments, unless stated otherwise. For SDXL, we set $L^{\mathrm{feat}}=\{0\}_{\mathrm{decoder}}$, $L^{\mathrm{self}}=\{0,1,2\}_{\mathrm{decoder}}$, $L^{\mathrm{app}}=\{1,2,3,4\}_{\mathrm{decoder}}\cup\{2,3,4,5\}_{\mathrm{encoder}}$, and $\tau^{\mathrm{s}}=\tau^{\mathrm{a}}=0.6$. We sample $\mathbf{I}^{\mathrm{o}}$ with $50$ steps of DDIM sampling and set $\eta=1$ [33], doing self-recurrence for $n^{\mathrm{r}}=2$ for $\tau^{\mathrm{r}}_{0}=0.1$ and $\tau^{\mathrm{r}}_{1}=0.5$. We implement Ctrl-X with Diffusers [37] and run all experiments on a single NVIDIA A6000 GPU.
Appendix B Extension to prompt-driven controllable generation

### v2

@@ Abstract
Recent controllable generation approaches such as FreeControl [24] and Diffusion Self-Guidance [7] bring fine-grained spatial and appearance control to text-to-image (T2I) diffusion models without training auxiliary modules. However, these methods optimize the latent embedding for each type of score function with longer diffusion steps, making the generation process time-consuming and limiting their flexibility and use. This work presents Ctrl-X, a simple framework for T2I diffusion controlling structure and appearance without additional training or guidance. Ctrl-X designs feed-forward structure control to enable the structure alignment with a structure image and semantic-aware appearance transfer to facilitate the appearance transfer from a user-input image. Extensive qualitative and quantitative experiments illustrate the superior performance of Ctrl-X on various condition inputs and model checkpoints. In particular, Ctrl-X supports novel structure and appearance control with arbitrary condition images of any modality, exhibits superior image quality and appearance transfer compared to existing works, and provides instant plug-and-play functionality to any T2I and text-to-video (T2V) diffusion model.
@@ 1 Introduction
To tackle these limitations, we present Ctrl-X, a simple training-free and guidance-free framework for T2I diffusion with structure and appearance control. We name our method “Ctrl-X” because we reformulate the controllable generation problem by ‘cutting’ (and ‘pasting’) two tasks together: spatial structure preservation and semantic-aware stylization. Our insight is that diffusion feature maps capture rich spatial structure and high-level appearance from early diffusion steps sufficient for structure and appearance control without guidance. To this end, Ctrl-X employs feature injection and spatially-aware normalization in the attention layers to facilitate structure and appearance alignment with user-provided images. By being guidance-free, Ctrl-X eliminates additional optimization overhead and sampling steps, resulting in a $35$-fold increase in inference speed compared to guidance-based methods. Figure 1 shows sample generation results. Moreover, Ctrl-X supports arbitrary structure conditions beyond natural images and can be applied to any T2I and even text-to-video (T2V) diffusion models. Extensive quantitative and qualitative experiments, along with a user study, demonstrate the superior image quality and appearance alignment of our method over prior works.
@@ 4.1 Feed-forward structure control
In contrast to [36, 18, 24], we do not perform inversion and instead directly use forward diffusion (Equation 1) to obtain $\mathbf{x}^{\mathrm{s}}_{t}$. We observe that $\mathbf{x}^{\mathrm{s}}_{t}$ obtained via the forward diffusion process contains sufficient structure information even at very early/high time steps, as shown in Figure 2. This also reduces appearance leakage common to inversion-based methods observed by FreeControl [24]. We study our feed-forward structure control method in Sections 5.1 and 5.2.
@@ 5.1 T2I diffusion with structure and appearance control
Baselines.    For training-based methods, ControlNet [44] and T2I-Adapter [25] learn an auxiliary module that injects a condition image into a pretrained diffusion model for structure alignment. We then combine them with IP-Adapter [43], a trained module for image prompting and thus appearance transfer. Uni-ControlNet [46] adds a feature extractor to ControlNet to achieve multi-image structure control of selected condition types, along with image prompting for global/appearance control. Splicing ViT Features [35] trains a U-Net from scratch per source-appearance image pair to minimize their DINO-ViT self-similarity distance and global [CLS] token loss. (For structure conditions not supported by a training-based baseline, we convert them to canny edge maps.) For guidance-based methods, FreeControl [24] enforce structure and appearance alignment via backpropagated score functions computed from diffusion feature subspaces. For guidance-free methods, Cross-Image Attention [1] manipulates attention weights to transfer appearance while maintaining structure. We run all methods on SDXL v1.0 [27] when possible and on their default base models otherwise.
@@
Evaluation metrics.    For quantitative evaluation, we report two widely-adopted metrics: DINO Self-sim measures the self-similarity distance [35] between the structure and output image in the DINO-ViT [6] feature space, where a lower distance indicates better structure preservation; DINO-I measures the cosine similarity between the DINO-ViT [CLS] tokens of the appearance and output images [30], where a higher score indicates better appearance transfer.
@@
Qualitative results.    As shown in Figures 4 and 5, Ctrl-X faithfully preserves structure from structure images ranging from natural images and ControlNet-supported conditions (e.g. HED, segmentation) to in-the-wild conditions (e.g. wireframe, 3D mesh) not possible in prior training-based methods while adeptly transferring appearance from the appearance image with semantic correspondence. Moreover, as shown in Figure 6, Ctrl-X is capable of multi-subject generation, capturing strong semantic correspondence between different subjects and the background, achieving balanced structure and appearance alignment. On the contrary, ControlNet + IP-Adapter [44, 43] often fails to maintain the structure and/or transfer the subjects’ or background’s appearances.
Comparison to baselines.    Figure 5 and Table 2 compare Ctrl-X to the baselines for qualitative and quantitative results, respectively. Moreover, our user study in Table 4, Appendix A shows the human preference percentages of how often participants preferred Ctrl-X over each of the baselines on result quality, structure fidelity, appearance fidelity, and overall fidelity.
@@
For training-based and guidance-based methods, despite Uni-ControlNet [46] and FreeControl’s [24] stronger structure preservation (smaller DINO self-similarity), they generally struggle to enforce faithful appearance transfer and yield worse DINO-I scores, which is particularly visible in Figure 5 row 1 and 3. Since the training-based methods combine a structure control module (ControlNet [44] and T2I-Adapter [25]) with a separately-trained appearance transfer module IP-Adapter [43], the two modules sometimes exert conflicting control signals at the cost of appearance transfer (e.g. row 1)—and for ControlNet, structure preservation as well. For Uni-ControlNet, compressing the appearance image to a few prompt tokens results in often inaccurate appearance transfer (e.g. rows 4 and 5) and structure bleed artifacts (e.g. row 6). For FreeControl, its appearance score function from extracted embeddings may not sufficiently capture more complex appearance correspondences, which, along with needing per-image hyperparameter tuning, results in lower contrast outputs and sometimes failed appearance transfer (e.g. row 4). Moreover, despite Splicing ViT Features [35] having the best self-similarity and DINO-I scores in Table 2, Figure 5 reveals that its output images are often blurry while displaying structure image appearance leakage with non-natural images (e.g. row 3, 5, and 6). It benchmarks well because its per-image training minimizes DINO metrics directly.
There is a trade-off between structure consistency (self-similarity) and appearance similarity (DINO-I), as these are competing metrics—increasing structure preservation corresponds to worse appearance similarity, which we show in Figure 11, Appendix B by varying controls schedules. As single metrics are not representative of overall method performance, we survey overall fidelity in our user study (Table 4, Appendix A), where Ctrl-X achieved best overall fidelity while matching result quality, structure fidelity, and appearance fidelity with training-based methods, showcasing our method’s ability to balance the conflicting, disentangled tasks of structure and appearance control.
@@
Guidance-free baseline Cross-Image Attention [1], in contrast, is less robust and more sensitive to the structure image, as the inverted structure latents contain strong appearance information. This causes both poorer structure alignment and frequent appearance leakage or artifacts (e.g. row 6) from the structure to the output images, resulting in worse DINO self-similarity and DINO-I scores. Similarly, Ctrl-X results are consistently preferred over Cross-Image Attention ones in our user study across all metrics (Table 4, Appendix A). In practice, we find Cross-Image Attention to be sensitive to its domain name, which is used for attention masking to isolate subjects, and it thus sometimes fails to produce outputs with cross-modal pairs (e.g. wireframes to photos).
Figure 6: Multi-subject generation.    Ctrl-X is capable of multi-subject generation with semantic correspondence between appearance and structure images across both subjects and backgrounds. In comparison, ControlNet + IP-Adapter [44, 43] often fails at transferring all subject appearances.
[…]
Method Training Preprocessing time (s) Inference latency (s) Total time (s) Peak GPU memory usage (GiB) Splicing ViT Features [35] 0.00 1557.09 1557.09 3.95 Uni-ControlNet [46] 0.00 6.96 6.96 7.36 ControlNet + IP Adapter [44, 43] 0.00 6.21 6.21 18.09 T2I-Adapter + IP-Adapter [25, 43] 0.00 4.37 4.37 13.28 Cross-Image Attention [1] 18.33 24.47 42.80 8.85 FreeControl [24] 239.36 139.53 378.89 44.34 Ctrl-X (ours) 0.00 10.91 10.91 11.51
@@
Inference efficiency.    We study the inference time, preprocessing time, and peak GPU memory usage of our method compared to the baselines, all with base model SDXL v1.0 except Uni-ControlNet (SD v1.5), Cross-Image Attention (SD v1.5), and Splicing ViT Features (U-Net). Table 1 reports the average inference time using a single NVIDIA H100 GPU. Ctrl-X is slightly slower than training-based ControlNet (1.76×1.76$\times$1.76 ×) and T2I-Adapter (2.50×2.50$\times$2.50 ×) with IP-Adapter yet significantly faster than per-image-trained Splicing ViT (0.0070×0.0070$\times$0.0070 ×), guidance-based FreeControl (0.029×0.029$\times$0.029 ×), and guidance-free Cross-Image Attention (0.25×0.25$\times$0.25 ×). Moreover, for methods with SDXL v1.0 as the base model, Ctrl-X has lower peak GPU memory usage than training-based methods and significantly lower memory than training-free methods. Our training-free and guidance-free method achieves comparable run time and peak GPU memory usage compared to training-based methods, indicating its flexibility.
Table 2: Quantitative comparison of structure and appearance control.    Ctrl-X consistently outperforms both training-based and training-free methods in appearance alignment and shows comparable or better structure preservation compared to training-based and guidance-free methods, measured by DINO ViT self-similarity [35] and DINO-I [30], respectively.
Method Training Natural image ControlNet-supported New condition Self-sim $\downarrow$ DINO-I $\uparrow$ Self-sim $\downarrow$ DINO-I $\uparrow$ Self-sim $\downarrow$ DINO-I $\uparrow$ Splicing ViT Features [35] 0.030 0.907 0.043 0.864 0.037 0.866 Uni-ControlNet [46] 0.045 0.555 0.096 0.574 0.073 0.506 ControlNet + IP-Adapter [44, 43] 0.068 0.656 0.136 0.686 0.139 0.667 T2I-Adapter + IP-Adapter [25, 43] 0.055 0.603 0.118 0.586 0.109 0.566 Cross-Image Attention [1] 0.145 0.651 0.196 0.510 0.175 0.570 FreeControl [24] 0.058 0.572 0.101 0.585 0.089 0.567 Ctrl-X (ours) 0.057 0.686 0.121 0.698 0.109 0.676
Figure 7: Qualitative comparison of conditional generation.    Ctrl-X displays comparable structure control and superior prompt alignment to training-based methods, and it also has better image quality and is more robust than guidance-based and guidance-free methods across different conditions.
@@
Extension to prompt-driven conditional generation.    Ctrl-X also supports prompt-driven conditional generation, where it generates an output image complying with the given text prompt while aligning with the structure from the structure image, as shown in Figures 4 and 7. Inspired by FreeControl [24], instead of a given $\mathbf{I}^{\mathrm{a}}$, Ctrl-X can jointly generate $\mathbf{I}^{\mathrm{a}}$ based on the text prompt alongside $\mathbf{I}^{\mathrm{o}}$, where we obtain $\mathbf{x}^{\mathrm{a}}_{t-1}$ via denoising with Equation 2 from $\mathbf{x}^{\mathrm{a}}_{t}$ without control. Baselines, qualitative and quantitative analysis, and implementation details are available in Appendix C.
@@
Extension to video diffusion models.    Ctrl-X is training-free, guidance-free, and demonstrates competitive runtime. Thus, we can directly apply our method to text-to-video (T2V) models, as seen in Figure 17, Appendix D. Our method closely aligns the structure between the structure and output videos while transferring temporally consistent appearance from the appearance image.
@@ 5.2 Ablations
(a) Ablation on control (b) Ablation on appearance transfer method
Refer to caption Refer to caption
(c) Ablation on inversion vs. our method
Figure 8: Ablations.    We study ablations on control, appearance transfer method, and inversion.
@@
Appearance transfer method.    As we consider appearance transfer as a stylization task, we compare our appearance statistics transfer with and without attention weighting in Figure 8(b). Without weighting (equivalent to AdaIN [15]), we have global normalization which ignores the semantic correspondence between the appearance and output images, so the outputs are low-contrast.
@@ 6 Conclusion
We present Ctrl-X, a training-free and guidance-free framework for structure and appearance control of any T2I and T2V diffusion model. Ctrl-X utilizes pretrained T2I diffusion model feature correspondences, supports arbitrary structure image conditions, works with multiple model architectures, and achieves competitive structure preservation and superior appearance transfer compared to training- and guidance-based methods while enjoying the low overhead benefits of guidance-free methods. As shown in Figure 9, the key limitation of Ctrl-X is the semantic-aware appearance transfer method may fail to capture the target appearance when the instance is small because of the low resolution of the feature map. We hope our method and findings can unveil new possibilities and research on controllable generation as generative models become bigger and more capable.
@@
Acknowledgements.    This work was supported by the NSF Grants CCRI-2235012 and RI-2339769, the UCLA–Amazon Science Hub, and the Intel Rising Star Faculty Award.
Appendix A Method, implementation, and evaluation details
@@
$n^{\mathrm{r}}$ times for (normalized) time steps $t\in[\tau^{\mathrm{r}}_{0},\tau^{\mathrm{r}}_{1}]$, where $\tau^{\mathrm{r}}_{0},\tau^{\mathrm{r}}_{1}\in[0,1]$. Notably, the self-recurrence steps occur without structure nor appearance control, and we observe generally lower artifacts and slightly better appearance transfer when self-recurrence is enabled.
Comparison to prior works.    We compare Ctrl-X to prior works in terms of capabilities in Table 3. Compared to baselines, our method is the only work which supports appearance and structure control with any structure conditions, while being training-free and guidance-free.
@@
Experiment hyperparameters.    For both T2I diffusion with structure and appearance control and structure-only conditional generation, we use Stable Diffusion XL (SDXL) v1.0 [27] for all Ctrl-X experiments, unless stated otherwise. For SDXL, we set $L^{\mathrm{feat}}=\{0\}_{\mathrm{decoder}}$, $L^{\mathrm{self}}=\{0,1,2\}_{\mathrm{decoder}}$, $L^{\mathrm{app}}=\{1,2,3,4\}_{\mathrm{decoder}}\cup\{2,3,4,5\}_{\mathrm{encoder}}$, and $\tau^{\mathrm{s}}=\tau^{\mathrm{a}}=0.6$. We sample $\mathbf{I}^{\mathrm{o}}$ with $50$ steps of DDIM sampling and set $\eta=1$ [33], doing self-recurrence for $n^{\mathrm{r}}=2$ for $\tau^{\mathrm{r}}_{0}=0.1$ and $\tau^{\mathrm{r}}_{1}=0.5$. We implement Ctrl-X with Diffusers [37] and run all experiments on a single NVIDIA A6000 GPU, except evaluating inference efficiency in Table 1 where we run on a single NVIDIA H100 GPU.
More details on evaluation metrics.    To evaluate structure and appearance control results (Table 2), we report DINO Self-sim and DINO-I. For DINO Self-sim, we compute the self-similarity (i.e., mean squared error) between the structure and output image in the DINO-ViT [6] feature space, where we use the base-sized model with patch size $8$ following Splicing ViT Features [35]. For DINO-I, we compute the cosine similarity between the DINO-ViT [CLS] tokens of the appearance and output images, where we use the small-sized model with patch size $16$ following DreamBooth [30].
To evaluate prompt-driven controllable generation results (Table 5), we report DINO-Self-sim, CLIP score, and LPIPS. DINO Self-sim is computed the same way as structure and appearance control metrics. For CLIP score, we compute the cosine similarity between the output image and text prompt in the CLIP embedding space, where we use the large-sized model with patch size $14$ (ViT-L/14) following FreeControl [24]. For LPIPS, we compute the appearance deviation of the output image from the structure image, where we use the official lpips package [45] with AlexNet (net="alex").
Method Structure control Appearance control Training-free Guidance-free Natural images In-the-wild conditions Uni-ControlNet [46] ControlNet [44] (+ IP-Adapter [43]) T2I-Adapter [25] (+ IP-Adapter [43]) SDEdit [23] Prompt2Prompt [11] Plug-and-Play [36] InfEdit [41] Splicing ViT Attention [35] Cross-Image Attention [1] FreeControl [24] Ctrl-X (ours)
Table 4: Qualitative comparison of structure and appearance control via user study.    The human preference percentages here show how often the participants preferred Ctrl-X over each of the baselines on result quality, structure fidelity, appearance fidelity, and overall fidelity. Ctrl-X consistently outperforms training-free baselines and is competitive with training-based ones, especially with overall fidelity, showcasing Ctrl-X’s ability to balance structure and appearance control.
Method Training Result quality $\uparrow$ Structure fidelity $\uparrow$ Appearance fidelity $\uparrow$ Overall fidelity $\uparrow$ Splicing ViT Features [35] 95% 87% 56% 78% Uni-ControlNet [46] 86% 17% 96% 74% ControlNet + IP-Adapter [44, 43] 46% 61% 41% 50% T2I-Adapter + IP-Adapter [25, 43] 74% 53% 67% 58% Cross-Image Attention [1] 95% 83% 83% 83% FreeControl [24] 64% 48% 79% 74% Ctrl-X (ours) - - - -
User study.    We follow the setting of the user study from DenseDiffusion [17], where we compare Ctrl-X to baselines on structure and appearance control in Table 4, we display the average human preference percentages of how often participants preferred our method over each of the baselines. We randomly selected $15$ sample pairs from our dataset and then assigned each sample pair to $7$ methods: Splicing ViT Feature [35], Uni-ControlNet [46], ControlNet + IP-Adapter [44, 43], T2I-Adapter + IP-Adapter [25, 43], Cross-Image Attention [1], FreeControl [24], and Ctrl-X. We invited $10$ users to evaluate pairs of results, each consisting of our method, Ctrl-X, and a baseline method. For each comparison, users assessed $15$ pairs between Ctrl-X and each baseline, based on four criteria: “the quality of displayed images,” “the fidelity to the structure reference,” “the fidelity to the appearance reference,” and “overall fidelity to both structure and appearance reference,” which we denote result quality, structure fidelity, appearance fidelity, and overall fidelity, respectively. We collected $150$ comparison results for between Ctrl-X and each individual baseline method. We reported the human preference rate, which indicates the percentage of times participants preferred our results over the baselines. The user study demonstrates that Ctrl-X outperforms training-free baselines and has a competitive performance compared to training-based baselines.
The user study (Figure 10) is conducted via Amazon Mechanical Turk.
Figure 10: User study interface.    A screenshot of our user study’s interface.
Appendix B Structure and appearance schedules and higher-level conditions
Ctrl-X has two hyperparameters, structure control schedule ($\tau^{\mathrm{s}}$) and appearance control schedule ($\tau^{\mathrm{a}}$), which enable finer control over the influence of the structure and appearance images on the output. As structure alignment and appearance transfer are conflicting tasks, controlling the two schedules allows the user to determine the best tradeoff between the two. The default values of $\tau^{\mathrm{s}}=0.6$ and $\tau^{\mathrm{a}}=0.6$ we choose merely works well for most—but not all—structure-appearance image pairs. Particularly, this control enables better results for challenging structure-appearance pairs and allows our method to be used with higher-level conditions without clear subject outlines.
Figure 11: Ablation on control schedules.    By varying Ctrl-X’s structure and appearance control schedules ($\tau^{\mathrm{s}}$ and $\tau^{\mathrm{a}}$), we change the influence of the structure and appearance images on the output.
Effect of control schedules.    We vary structure and appearance control schedules ($\tau^{\mathrm{s}}$ and $\tau^{\mathrm{a}}$) as seen in Figure 11. Decreasing structure control can make cross-class structure-appearance pairs (e.g., horse normal map with puppy appearance) look more realistic, as doing so trades strict structure adherence for more sensible subject shapes in challenging scenarios. Decreasing appearance control trades appearance alignment for less artifacts. Note that, generally, $\tau^{\mathrm{s}}\leq\tau^{\mathrm{a}}$, as structure control requires appearance transfer to realize the structure information and avoid structure image appearance leakage, most prominently demonstrated in Figure 8(a).
Figure 12: Higher-level structure conditions.    By decreasing the structure schedule $\tau^{\mathrm{s}}$ (from the default $0.6$ to 0.30.30.30.30.50.50.50.5), Ctrl-X can handle higher-level structure conditions such as bounding boxes (left) and human pose skeletons/keypoints (right).
Higher-level structure conditions.    By decreasing the structure control schedule $\tau^{\mathrm{s}}$ from the default $0.6$ to 0.30.30.30.30.50.50.50.5, Ctrl-X can handle sparser and higher-level structure conditions such as bounding boxes and human post skeletons/keypoints, shown in Figure 12. Not only does this make our method applicable to other higher-level control types, it also generally reduces structure image appearance leakage with challenging structure conditions.
Appendix C Extension to prompt-driven controllable generation